OpenStudy (anonymous):

The sum of two numbers is 170 and their difference is 26. Find the two numbers. Write the solution as an ordered pair.

OpenStudy (anonymous):

x + y = 170 x - y = 26 ======= 2x = 196 x = 196/2 x = 98 Put the value of x in any equation 98 + y = 170 y = 170 - 98 y = 72 The two numbers are 98 and 72 Ordered pair: (98 , 72)

OpenStudy (nurali):

x + y = 170----(1) x - y = 26-----(2) Add equation 1 and 2 2x = 196 x = 196/2 x = 98 put x in equation (1) x + y = 170 98+y=170 y = 72 (x,y) =(98 , 72)

OpenStudy (anonymous):

Thank you both of you I have many question im not that good at math, are you ready for another question??:D

OpenStudy (anonymous):

,To skate at Roller Heaven, each person must pay a membership fee (m) and a fee for each session (f). Art attended five sessions and paid $25 in all. Later, when he had attended 10 sessions, he paid $40 in all. How much is the membership fee and how much are the sessions. Write a system of two equations with the variables m and f in each to represent this problem.

OpenStudy (anonymous):

Yeah sure why not :) Make different threads for different questions :) Thanks :)

OpenStudy (anonymous):

In the first part m+5f=25 And for the second 10f=40 f=40/10 f=4 Put f=4 into the 1st part, that is m+5f=25 m+5*4=25 m+20=25 m=5 So, sessions fee is 4 and membership fee is 5. I hope I'm clear :)

OpenStudy (anonymous):

thank you, you are really good at math!

OpenStudy (anonymous):

Your welcome :) And thanks :D